OpenStudy (chaylaceyx3):

Find the point on the terminal side of θ =-3pi/4 that has an x coordinate of -1

OpenStudy (anonymous):

can you draw a picture of what this is asking?

OpenStudy (anonymous):

what quadrant is \(\theta \) in?

OpenStudy (chaylaceyx3):

No, im supposed to use the unit cirle and stuff... really it would help me to know if -3pi/4 is the same as a positive 3pi/4. Because 3pi/4 is 135, but im not sure if a -3pi/4 would be 225?

OpenStudy (anonymous):

yes, it's basically the same but a negative angle is just rotating clockwise. so instead of rotating 3/4 pi counter-clockwise, a negative angle would rotate 3/4pi clockwise....

OpenStudy (chaylaceyx3):

so it would be 225?

OpenStudy (anonymous):

well 225 would be a coterminal angle and u can use that to answer this question...

OpenStudy (anonymous):

ie, 225 degrees is in the same quadrant as \(\large -\frac {3\pi}{4} \)

OpenStudy (chaylaceyx3):

Okay, I think that helps me! Thanks (: